# PreCi: Pre-training and Continual Improvement of Humanoid Locomotion via Model-Assumption-based Regularization

### v1

@@ VI-C1 Robustness Test
PreCi 100 $\mathbf{8.4}$ 100 $\mathbf{9.1}$ 100 $\mathbf{18.9}$
@@
We first evaluate robustness by commanding the robot to traverse uphill platforms with three different slopes and measure the success rate and linear velocity tracking error with respect to the commanded velocity, as shown in Table II.
Overall, PreCi, IFM, and FullReg demonstrate strong robustness, while MBC and PureRL exhibit significantly degraded performance. MBC consistently performs the worst, whereas controllers that imitate it (PreCi, FullReg, and IFM) show substantially improved robustness. This improvement can be attributed to training the policies on diverse terrains with reinforcement learning objectives, which improves their ability to handle challenging environments. Although PureRL successfully navigates the final slope terrain in the MuJoCo test, it achieves much lower success rates in AR-Sim. This highlights the importance of incorporating the periodic gait style of MBC, which facilitates transfer to different simulators and potentially to real-world environments.
Among the methods achieving a 100% success rate, PreCi outperforms the others in tracking accuracy, exhibiting the smallest increase in tracking error as the slope angle increases. In contrast, FullReg shows worse tracking performance, suggesting that full regularization with MBC degrades tracking accuracy in the presence of model assumption errors. We further analyze the effect of MAR in Section VI-E in the presence of model-assumption violation.
@@ VI-C2 Random Velocity Tracking Test
Figure 6: Mean undiscounted return of tracking rewards in the AR-Sim random velocity tracking test. The command range is wider than that used during training. Policies learned with regularization, PreCi and FullReg, show superior performance compared to those without MBC regularization.
@@
PreCi outperforms all baselines, achieving a slightly higher return than FullReg. In contrast, IFM shows worse tracking performance than the other regularization-based methods. This is mainly due to the robot frequently tripping over its own foot and falling when given sudden changes in lateral or angular commands.
Despite its robust performance in the training environment, PureRL demonstrates the worst performance among the baselines, performing even worse than MBC. This result aligns with the observations from the previous experiments, supporting the claim that motion distillation from MBC, whether through pre-training alone (IFM) or in combination with regularization (PreCi and FullReg), enhances the controller’s robustness to domain shifts.
@@ VI-D Hardware Experiments
We deploy the controllers on hardware across five different scenarios, both indoor and outdoor, as illustrated in Fig. 1. The results are summarized in Table III.
@@ VI-D1 Indoor Experiments
In indoor experiments, we measure the maximum forward velocity on flat terrain and the success rate of traversing the step-and-slip platform. This platform consists of a whiteboard covered with poppy seeds to induce slipping, and the robot is commanded to traverse it at 0.3 m/s. The success rate is measured over three trials.
PreCi achieves the best performance among all baselines, reaching a maximum velocity of 1.5 m/s while maintaining a 100% success rate in the step-and-slip test. In contrast, MBC and FullReg reach maximum velocities of only 0.5 m/s and 1.1 m/s, respectively. Both controllers fail the step-and-slip test in all trials, slipping in the middle of the whiteboard. Despite being trained with the same RL objective as PreCi, FullReg shows limited performance and exhibits a failure pattern similar to MBC.
One possible reason for PreCi’s performance gain is its ability to handle model assumption violations at higher velocities, whereas MBC requires further tuning of its parameters [44]. Additionally, by adjusting reliance on MBC action regularization in certain states and placing greater emphasis on the RL objective, PreCi better adapts to prevent slips.
IFM struggles with stability during the maximum forward velocity test, frequently tripping over its own foot. For safety, we halted the experiment at 1.0 m/s. PureRL performs the worst in hardware tests. Its maximum velocity could not be measured due to poor tracking performance, and it fails the step-and-slip test as its foot becomes stuck on the whiteboard surface. As observed in the AR-Sim test, PureRL does not transfer well to new environments beyond its training domain, making it unsuitable for sim-to-real transfer.
@@ VI-D2 Outdoor Experiments
Both PreCi and FullReg successfully traverse all terrain types, including deformable surfaces such as sand, which were not included during training. In contrast, MBC shows limited robustness, struggling even on mildly challenging terrain. Please refer to the supplemental video for details.
@@ VI-E Effect of MAR in Training and Testing
Figure 7: Left: an adaptive weight distribution against model assumption violation. Right: a histogram of regularization loss weight during PreCi training on both uphill and flat terrain. We observe more unreliable samples with low regularization weight (region 1) and fewer reliable samples with high weight (region 3) in the uphill case, reflecting how PreCi filters out unreliable regularization in violation-prone regions.
Figure 8: Relationship between model-assumption violations and linear velocity tracking errors during a 14-degree uphill test. FullReg exhibits higher tracking errors than PreCi when model-assumption violations are more pronounced, while both methods show similar low error distributions when violations are less significant.
Although FullReg demonstrates robust performance in both simulation and hardware, its capabilities are often limited compared to PreCi when model-assumption violations occur, including the maximum velocity it can reach and tracking ability. We analyze how MAR takes effect when there are model-assumption violations in training and testing.
@@ VI-E1 Analysis of MAR at Training
Overall, MAR successfully detects model-assumption violations and leverages them to adjust the regularization weights dynamically. On flat terrain, violations are infrequent except under high-velocity commands. In this case, a large proportion of MBC actions remain reliable (region 3 in Fig. 7), resulting in a higher regularization weight. In contrast, more violations are observed on sloped terrain (region 1 in Fig. 7), leading to lower weights. This enables PreCi to downweight suboptimal regularization, thereby improving performance in both simulated and real-world environments.
@@ VI-E2 Analysis of MAR at Testing
We examine the relationship between model-assumption violations and linear velocity errors for PreCi and FullReg to evaluate the effectiveness of the MAR during testing on various terrains. Fig. 8 shows the relationship between model-assumption violation (Z-velocity) and linear velocity error in the 14-degree uphill test. A linear model is fitted to quantify the correlation.
During testing, the model trained with MAR demonstrates consistent performance even in the presence of model-assumption violations. For FullReg, linear velocity tracking error increases as violations become more frequent with a correlation of 0.4110. In contrast, PreCi maintains a more robust tracking performance with a correlation of 0.1562.
Interestingly, MAR improves not only performance on challenging terrains but also on simpler ones. While PreCi performs better on uneven terrain and slopes in MuJoCo simulations, it also outperforms other methods in the Maximum Forward Velocity test on flat terrain and the Step-and-Slip test on hardware, as shown in Table III. This improvement may result from increased model-assumption violations caused by high-speed or unstable locomotion even in simple scenarios, or it may reflect PreCi’s ability to achieve more general performance by ignoring irrelevant samples during training.
@@ VII CONCLUSIONS
We propose a novel learning framework, PreCi, that combines model-based control and learning-based approaches to effectively train a robust humanoid locomotion control policy. PreCi integrates several components, such as pre-training through imitation of a model-based controller, fine-tuning via reinforcement learning, and model assumption-based regularization (MAR), to enhance the robustness and adaptability of humanoid robots across diverse and challenging tasks. Our approach addresses key limitations of prior methods, such as the limited performance of model-based controllers and the catastrophic forgetting in IFM [23].
Through extensive simulation tests and hardware experiments on the Digit humanoid robot, we demonstrate that PreCi outperforms baseline methods in terrain robustness, velocity tracking, and sim-to-real transfer. PreCi achieves a maximum forward velocity of 1.5 m/s and can navigate complex terrains, including slippery surfaces, slopes, uneven ground, and deformable sand, with zero-shot deployment. The incorporation of MAR enables the policy to adapt to scenarios where model assumptions are violated, resulting in superior performance compared to fully regularized methods.
One future extension of this work is to apply MAR to a broader range of model-based assumptions. This includes an in-depth investigation into methods for quantifying model-assumption violations in the single rigid body model. In addition, another natural extension is to adapt the framework for more dynamic and high-impact motions, such as running or jumping. This could involve refining the model assumptions or regularization techniques to better handle faster dynamics and increased instability, broadening the framework’s applicability to agile robotic behaviors.

### v2

@@ VI-C1 Robustness Test
PPF 100 $\mathbf{8.4}$ 100 $\mathbf{9.1}$ 100 $\mathbf{18.9}$
@@
We first evaluate robustness by commanding the robot to traverse uphill platforms with three different slopes and measure the success rate and linear velocity tracking error with respect to the commanded velocity, as shown in Table II. A trajectory is considered successful if the robot traverses the 5-meter uphill section without falling.
Overall, PPF, IFM, and FullReg demonstrate strong robustness, while MBC and PureRL exhibit significantly degraded performance. MBC consistently performs the worst, whereas controllers that imitate it (PPF, FullReg, and IFM) show substantially improved robustness. This improvement can be attributed to training the policies on diverse terrains with reinforcement learning objectives, which improves their ability to handle challenging environments. Although PureRL successfully navigates the final slope terrain in the MuJoCo test, it achieves much lower success rates in AR-Sim. This highlights the importance of incorporating the periodic gait style of MBC, which facilitates transfer to different simulators and potentially to real-world environments.
Among the methods achieving a 100% success rate, PPF outperforms the others in tracking accuracy, exhibiting the smallest increase in tracking error as the slope angle increases. In contrast, FullReg shows worse tracking performance, suggesting that full regularization with MBC degrades tracking accuracy in the presence of model assumption errors. We further analyze the effect of MAR in Section VI-E in the presence of model-assumption violation.
@@ VI-C2 Random Velocity Tracking Test
Figure 6: Mean undiscounted return of tracking rewards in the AR-Sim random velocity tracking test. The command range is wider than that used during training. Policies learned with regularization, PPF and FullReg, show superior performance compared to those without MBC regularization.
@@
PPF outperforms all baselines, achieving a slightly higher return than FullReg. In contrast, IFM shows worse tracking performance than the other regularization-based methods. This is mainly due to the robot frequently tripping over its own foot and falling when given sudden changes in lateral or angular commands.
Despite its robust performance in the training environment, PureRL demonstrates the worst performance among the baselines, performing even worse than MBC. This result aligns with the observations from the previous experiments, supporting the claim that motion distillation from MBC, whether through pre-training alone (IFM) or in combination with regularization (PPF and FullReg), enhances the controller’s robustness to domain shifts.
@@ VI-D Hardware Experiments
We deploy the controllers on hardware across six different scenarios, both indoor and outdoor, as illustrated in Fig. 1. The results are summarized in Table III.
@@ VI-D1 Indoor Experiments
In indoor experiments, we measure the maximum forward velocity on flat terrain and the success rate of traversing the step-and-slip platform. This platform consists of a whiteboard covered with poppy seeds or olive oil to induce slipping, and the robot is commanded to traverse it at 0.3 m/s. The success rate is measured over five trials.
PPF achieves the best performance among all baselines, reaching a maximum velocity of 1.5 m/s while maintaining a 100% success rate in the step-and-slip test. In contrast, MBC and FullReg reach maximum velocities of only 0.5 m/s and 1.1 m/s, respectively. Both controllers fail the step-and-slip test in all trials, slipping in the middle of the whiteboard. Despite being trained with the same RL objective as PPF, FullReg shows limited performance and exhibits a failure pattern similar to MBC.
One possible reason for PPF’s performance gain is its ability to handle model assumption violations at higher velocities, whereas MBC requires further tuning of its parameters [26]. Additionally, by adjusting reliance on MBC action regularization in certain states and placing greater emphasis on the RL objective, PPF better adapts to prevent slips.
IFM struggles with stability during the maximum forward velocity test, frequently tripping over its own foot. For safety, we halted the experiment at 1.0 m/s. PureRL performs the worst in hardware tests. Its maximum velocity could not be measured due to poor tracking performance, and it fails the step-and-slip test as its foot becomes stuck on the whiteboard surface. PureRL struggles to transfer to new domains beyond its training domain unless the motion is carefully designed through extensive reward engineering or it is trained with additional components for sim-to-real, as demonstrated in [9].
@@ VI-D2 Outdoor Experiments
Both PPF and FullReg successfully traverse all terrain types, including deformable surfaces such as sand, which were not included during training. In contrast, MBC shows limited robustness, struggling even on mildly challenging terrain. Please refer to the supplemental video for details.
@@ VI-E Effect of MAR in Training and Testing
Figure 7: Left: an adaptive weight distribution against model assumption violation. Right: a histogram of regularization loss weight during PPF training on both uphill and flat terrain. We observe more unreliable samples with low regularization weight (region 1) and fewer reliable samples with high weight (region 3) in the uphill case, reflecting how PPF filters out unreliable regularization in violation-prone regions.
Figure 8: Relationship between model-assumption violations and linear velocity tracking errors during a 14-degree uphill test. FullReg exhibits higher tracking errors than PPF when model-assumption violations are more pronounced.
Although FullReg demonstrates robust performance in both simulation and hardware, its capabilities are often limited compared to PPF when model-assumption violations occur, including the maximum velocity it can reach and tracking ability. We analyze how MAR takes effect when there are model-assumption violations in training and testing.
@@ VI-E1 Analysis of MAR at Training
Overall, MAR successfully detects model-assumption violations and leverages them to adjust the regularization weights dynamically. On flat terrain, violations are infrequent except under high-velocity commands. In this case, a large proportion of MBC actions remain reliable (region 3 in Fig. 7), resulting in a higher regularization weight. In contrast, more violations are observed on sloped terrain (region 1 in Fig. 7), leading to lower weights. This enables PPF to downweight suboptimal regularization, thereby improving performance in both simulated and real-world environments.
@@ VI-E2 Analysis of MAR at Testing
We examine the relationship between model-assumption violations and linear velocity errors for PPF and FullReg to evaluate the effectiveness of the MAR during testing. Fig. 8 shows the relationship between model-assumption violation (Z-velocity) and linear velocity error in the 14-degree uphill test.
During testing, the model trained with MAR demonstrates consistent performance even in the presence of model-assumption violations. For FullReg, linear velocity tracking error increases as violations become more frequent.
Interestingly, MAR improves not only performance on challenging terrains but also on simpler ones. While PPF performs better on uneven terrain and slopes in MuJoCo simulations, it also outperforms other methods in the Maximum Forward Velocity test on flat terrain and the Step-and-Slip test on hardware, as shown in Table III. This improvement may result from increased model-assumption violations caused by high-speed or unstable locomotion even in simple scenarios, or it may reflect PPF’s ability to achieve more general performance by ignoring irrelevant samples during training.
@@ VII CONCLUSIONS
We propose a novel learning framework, PPF, that combines model-based control and learning-based approaches to effectively train a robust humanoid locomotion control policy. PPF integrates several components, such as pre-training through imitation of a model-based controller, fine-tuning via reinforcement learning, and model assumption-based regularization (MAR), to enhance the robustness and adaptability of humanoid robots across diverse and challenging tasks. Our approach addresses key limitations of prior methods, such as the limited performance of model-based controllers and the catastrophic forgetting in IFM [23].
Through extensive simulation tests and hardware experiments on the Digit humanoid robot, we demonstrate that PPF outperforms baseline methods in terrain robustness, velocity tracking, and sim-to-real transfer. PPF achieves a maximum forward velocity of 1.5 m/s and can navigate complex terrains, including slippery surfaces, slopes, uneven ground, and deformable sand, with zero-shot deployment. The incorporation of MAR enables the policy to adapt to scenarios where model assumptions are violated, resulting in superior performance compared to fully regularized methods.
One future extension of this work is to apply MAR to a broader range of model-based assumptions. For example, MAR can utilize estimated foot slipperiness through stance foot velocity, as most high fidelity models assume zero contact velocity. Furthermore, applying MAR to more sophisticated model-based controllers could provide insights into whether incorporating stronger prior knowledge during pre-training enhances training efficiency and final performance. Another possible extension is to adapt the framework for more dynamic and high-impact motions, such as running or jumping. This could involve refining the model assumptions or regularization techniques to better handle faster dynamics and increased instability, broadening the framework’s applicability to agile robotic behaviors.